lative 4.3 Find the mean of the random variables described by each of the following cumu distribution functions 0 1,2 0yS2, w< 0, w> 10 (b)/Fx(x)=[1-exp(-2x)]u(x) ; 0, 4 1, y> 2; (d) Fz(z)=[1-exp(--2)]u(z).
Answer #1

We know that   E(X) =\int_{0}^{\infty} [1-F(x)]dx

a) So

10 2、10 w-

b)

E(y) =\int _{0}^{2}[1-{y^2\over 4}]dy= \left ( y-{y^3\over 12} \right )_{0}^{2}={5\over 3}

c)

2 2 0 0

d)

E(Z)